OpenStudy (anonymous):

The reaction for the combustion of octane (C8H18) is given in this reaction: 2C8H18 + 25O2 → 16CO2 + 18H2O How many grams of H2O are produced if 228 grams of octane are burned? 181 grams 228 grams 81 grams 324 grams

imqwerty (imqwerty):

\(2C_8H_{18}~+~25O_2\) ------->\(16CO_2~+~18H_2O\) here we see that 2 moles of \(C_8H_{18}\) (Octane) react to produce 18 moles of \(H_2O\) so 228 grams of Octane are burned moles of octane burned= \(\large \frac{mass}{molar~mass}\) so to find moles we gotta know the molar mass molar mass of Octane (\(C_8H_{18}\))= \(8 \times Molar~mass~of~C+18 \times Molar~mass~of~H\) molar mass of C is 12 and of H is 1 so we get \(Molar~mass~of~C_8H_{18}=114\) nw we have molar mass of octane and also the mass burned so moles of octane burned= \(\large \frac{mass}{molar~mass}\) moles of octane burned= \(\large \frac{228}{114}\) moles of octane burned= \(2\) moles we know that 2 moles of octane give 18 moles of \(H_2O\) so we get 18 moes of \(H_2O\) mass of 18 moles of \(H_2O\) =\(moles \times Molar~mass\) molar mass of \(H_2O\) is 18 so put it in mass of 18 moles of \(H_2O\) =\(18\times 18\) we get mass of \(H_2O\) used =324grams